An empty tank is filled by an amount of kerosene of mass 217.8 ton by using a hose where the kerosene emerges from its nozzle with a velocity of 7.7 m/s so the tank is filled during 1000s , calculate the radius of the hose nozzle.
(Knowing that : density of kerosene = 900 kg/m³
1
Expert's answer
2019-12-23T11:28:16-0500
"\frac{\Delta V}{\Delta t}=\rho Av=\pi r^2\rho v"

"\frac{217800}{1000}=\pi (900)(7.7)r^2"

"r=0.1\ m"